# On Pretraining Data Diversity for Self-Supervised Learning

### v1

@@ Abstract
We explore the impact of training with more diverse datasets, characterized by the number of unique samples, on the performance of self-supervised learning (SSL) under a fixed computational budget. Our findings consistently demonstrate that increasing pretraining data diversity enhances SSL performance, albeit only when the distribution distance to the downstream data is minimal. Notably, even with an exceptionally large pretraining data diversity achieved through methods like web crawling or diffusion-generated data, among other ways, the distribution shift remains a challenge. Our experiments are comprehensive with seven SSL methods using large-scale datasets such as ImageNet and YFCC100M amounting to over 200 GPU days. Code and trained models will be available at https://github.com/hammoudhasan/DiversitySSL.
@@
Figure 1: Impact of Diversity on Pretraining: Self-supervised learning (SSL) can be used to pretrain vision models on smaller datasets closely aligned to downstream task data, e.g., pets classification, hence with a small distribution shift (top, wild animals pretraining). Conversely, we could pretrain on an extensively varied dataset, with wide distribution differences (outdoor scenes, bottom). We demistify the role of pretraining diversity in SSL under a fixed computational budget, and highlight its effects in relationship to the distribution shift.
@@ 1 Introduction
Ultimately, our work provides a comprehensive analysis, from experiments worth 200 GPU days, on the interaction between data diversity and computational resources, and their impact on the performance of SSL models, with an aim to improve pretraining practices. Now, we will analyze the relevant literature.
@@ 2 Related Work
Pretraining at Scale   SSL is most effective when pretraining is conducted at scale, benefiting from large datasets and great computational resources. Initial attempts at large-scale pretraining were made via combining contrastive learning and clustering [55, 7]. The SEER model [25] was trained on 1 billion internet images using SwAV [8] applied on a RegNet [48] backbone. The importance of model scaling to leverage large-scale datasets is also shown in  [64, 27, 26], as well as the need for increasing training duration [64]. Additional strategies are necessary to achieve best performance at scale [66, 17, 23]. All considered works focus on reaching the best representations, without many considerations about training costs, thus encouraging unfair comparisons about the efficacy of data. A preliminary work [15] found for SimCLR [10] that increasing data from 500 thousand to 1 million images leads to a modest boost in downstream performance, but without limiting the training budget. While large uncurated datasets have been further explored [55], their efficacy in relationship with the distribution shift under normalized computation is still limitedly investigated.
Distribution Shift in SSL   Some studies have investigated how the pretraining data domain affects downstream performance on other domains. For various pretraining datasets, preliminary works [36, 15, 27, 40] observed that the best performing representations were learned on datasets that were similar to the downstream test task. Additionally, combining datasets before pretraining or combining self-supervised features learned from various datasets did not lead to significant improvements in [15]. In  [68, 51], they showcase higher generalization capabilities of SSL models compared to their supervised counterparts, for several downstream tasks in the presence of a distribution shift. In [59], they pretrained on several datasets, and observed different generalization depending on the object-centric or scene-centric appearance of the pretraining dataset. Furthermore, initial considerations on the impact of external data on downstream tasks under distribution shift have been proposed in [60]. Some compensate the distribution shift with the scale of training datasets [29]. Although partial information can be inferred by these works, there is still a lack of a fair, computation-normalized evaluation that allows to study the effects of the distribution shift in a controlled environment.
@@ 3 Preliminaries
where $\mathcal{L}_{\text{SSL}}$ is a loss function specific to the chosen SSL pretraining method.
After pretraining, the feature extractor $f_{\theta_{f}^{*}}$ can be deployed for various downstream tasks such as image classification, object detection, or segmentation. This is typically achieved by training a task-specific head. Alternatively, the feature extractor $f_{\theta_{f}^{*}}$ can either be fine-tuned or used together with a $k$-nearest neighbors (kNN) classifier.
Linear Probing   There are several ways to evaluate the performance of a self-supervised learning method such as linear probing [28, 14, 10], kNN [62, 71, 9, 8], and few-shot evaluation [25, 22]. Consistent with the general protocol established in the literature [10, 31, 67, 15], we use linear probing to measure the quality of the features extracted for classification tasks. The procedure is as follows: (1) a labeled downstream dataset, $\mathbb{D}_{\text{task}}$, consisting of image-class pairs $(\mathbf{x},\mathbf{y})\sim\mathbb{D}_{\text{task}}$, is selected for evaluation. (2) For each image $\mathbf{x}$, its representation is extracted using the pretrained feature extractor $f_{\theta_{f}^{*}}$, after which the linear classification head $t_{\theta_{t}}$, parameterized by $\theta_{t}$, is then applied to obtain $t_{\theta_{t}}(f_{\theta_{f}^{*}}(\mathbf{x}))$. (3) The linear head $t_{\theta_{t}}$ is optimized as follows:
@@
Note that only the parameters of the linear head $\theta_{t}$ are optimized, while the parameters $\theta^{*}_{f}$ of the feature extractor are kept frozen. The quality of features extracted by $f_{\theta_{f}^{*}}$ is directly inferred from the classification accuracy achieved on the test set of $\mathbb{D}_{\text{task}}$, which serves as the primary indicator of the quality of the extracted features.
@@ 4 Normalized Evaluation
$N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
(×$(\times 10^{3})$ (×$(\times 10^{-3})$ SimCLR B.T. BYOL SwAV VICReg MoCoV3 DINO
5 0.1
37.95
44.30
41.14
14.06
44.06
12.51
15.74
25 0.5
50.13
54.29
51.01
45.44
50.37
51.58
35.89
50 1.0 58.59 58.42 58.28 56.37 55.83 55.61 40.93
@@
$N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
(×$(\times 10^{3})$ (×$(\times 10^{-3})$ SimCLR B.T. BYOL SwAV VICReg MoCoV3 DINO
10 0.2
36.91
40.98
35.62
34.23
37.58
36.56
34.54
50 1.0
48.77
52.01
48.05
43.63
48.75
46.45
44.39
100 2.0 49.83 55.60 50.29 47.48 54.10 48.58 47.32
@@ 5.1 Performance on the Same Distribution
Insight (1)
When the distributions of the upstream and downstream tasks are the same, i.e., $\mathbb{D}_{\text{SSL}}$ = $\mathbb{D}_{\text{task}}$, in a normalized computation setup, increasing pretraining diversity $\mathcal{D}$ proves to be an effective strategy for enhancing SSL pretraining performance.
@@ 5.2 Increasing Data Diversity
As observed in Section 5.1, if $\mathbb{D}_{\text{SSL}}=\mathbb{D}_{\text{task}}$ having a more diverse pretraining set benefits SSL methods, under a normalized computation assumption. However, to increase diversity, sampling from the same distribution $\mathbb{D}_{\text{task}}$ to extend $\mathbb{D}_{\text{SSL}}$ is not always attainable. In fact, pretraining data is often sampled from a distribution different than $\mathbb{D}_{\text{task}}$. The added samples will then introduce a distribution shift between $\mathbb{D}_{\text{SSL}}$ and $\mathbb{D}_{\text{task}}$.
@@
Figure 2: Data Collection Strategies: We analyze strategies for collecting additional data ($\mathbb{A}$), i.e., collecting more source data, crawling the web or using synthetic images. Using a class prior (top row) simulates In-distribution trainings. We also collect images without class prior (bottom row) to analize the interactions between diversity and Out-of-distribution classes.
Setup   We use ImageNet-100 [56] as $\mathbb{D}_{\text{task}}$, and we construct multiple $\mathbb{D}_{\text{SSL}}$ to evaluate the effects of different data collection strategies and the distribution shift. We first introduce a set $\mathbb{B}$ composed of $65\times 10^{3}$ images from ImageNet-100 (50% of the dataset), which we use as a baseline for $\mathbb{D}_{\text{SSL}}$ with minimum diversity. We denote the 100 classes in ImageNet-100 as $\mathcal{T}_{\text{100}}$, and sample $\mathbb{B}$ uniformly including images from all classes. Next, we compare with pretraining on more diverse datasets as $\mathbb{D}_{\text{SSL}}$, imposing $\mathbb{D}_{\text{SSL}}=\mathbb{B}\cup\mathbb{A}$ where $\mathbb{A}$ includes $65\times 10^{3}$ images sampled with one of three strategies. To highlight the effects of the distribution shift, we include in $\mathbb{A}$ either images from In-distribution classes, i.e., selecting images from classes overlapping with $\mathcal{T}_{\text{100}}$, or images from Out-of-distribution classes, which do not overlap with $\mathcal{T}_{\text{100}}$. This is to study the effects of the distribution shift, since we do not assume access to downstream classes in real scenarios. For the Out-of-distribution samples, we define $\mathbb{A}$ as (1) random images sampled from the full ImageNet [18] dataset; (2) images crawled from Flickr, Bing, and DuckDuckGo; or (3) images generated with Stable Diffusion V2.1 [50]. We respectively call these sets $\mathbb{A}_{\text{Source}}^{\text{{{Out}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{Out}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{Out}}}}$. We also define their In-distribution counterparts as $\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{In}}}}$, respectively. Note how $\mathbb{B}\cup{\color[rgb]{0.87890625,0.4375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.87890625,0.4375,0}\pgfsys@color@rgb@stroke{0.87890625}{0.4375}{0}\pgfsys@color@rgb@fill{0.87890625}{0.4375}{0}\mathbb{A}_{\text{Source}}^{\text{{{In}}}}}$ is the full ImageNet-100, coherently with Section 5.1. Figure 2 shows each collection strategy, for which we provide implementation details in supplementary material. Although many factors (such as the appearance of images) impact the distribution shift, using a class prior imposes that any strategy using it would still result in closer distribution with respect to the same strategy without class priors [15]. We pretrain a ResNet-18 from scratch, with the same settings of Section 5.1, and $\mathcal{C}=50\times 10^{6}$. Note that this results in $\mathcal{D}=1.25\times 10^{-3}$ for pretraining on $\mathbb{B}$, and $\mathcal{D}=2.5\times 10^{-3}$ for pretraining on $\mathbb{B}\cup\mathbb{A}$, introducing a $\mathcal{D}$ difference of a factor of 2.
Results   We report the linear probing accuracies for pretraining on each $\mathbb{B}\cup\mathbb{A}$ as bars in Figure 3, showing the $\mathcal{C}$-normalized training on $\mathbb{B}$ as a dashed line. Surprisingly, without class priors (Out), including $\mathbb{A}_{\text{Source}}^{\text{{{Out}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{Out}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{Out}}}}$ underperforms compared to pretraining on $\mathbb{B}$ only. For instance, for SimCLR and while $\mathbb{B}$ scores 72.3% accuracy, increasing diversity reduces the accuracy by 1% in all cases. This might appear to conflict with our findings in Section 5.1, however, the inclusion of Out samples leads to $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$, since we sample only classes not included in $\mathbb{D}_{\text{task}}$. We infer that, with normalized $\mathcal{C}$, increasing $\mathcal{D}$ without distribution priors may negatively impact performance. Conversely, when class priors are available (In), increasing pretraining diversity $\mathcal{D}$ improves performance compared to $\mathbb{B}$ pretraining. For instance, pretraining on $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$ performs comparably to augmenting with additional ImageNet samples ($\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$), as in the case of SimCLR where the inclusion of $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$ scores only 0.2% lower than $\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$. Including $\mathbb{A}_{\text{Synthetic}}^{\text{{{In}}}}$ data also helps, Synthetic $\mathbb{A}_{\text{Synthetic}}^{\text{{{In}}}}$ data helps, although more limitedly due to the visual discrepancies between generated and real images. Ultimately, the effectiveness of diversity is linked to the distribution shift. These findings highlight the impact of the distribution shift on computationally-normalized SSL pretraining and help define evaluation practices for this task (see Section 7).
Insight (2)
When the distributions of the upstream and downstream tasks differ, $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$, and in a normalized computation setup, increasing pretraining diversity $\mathcal{D}$ may harm SSL pretraining performance, reflecting the adverse effects of distribution shift.
Table 2: Performance of ImageNet and YFCC100M SSL Pretraining on Various Downstream Tasks: We train ResNet-50 and ViT-B/16 under normalized computation ($\mathcal{C}=98\times 10^{6}$) using SimCLR (left) and MoCoV3 (right) on ImageNet and YFCC100M (YFCC) with multiple $\mathcal{D}$. Despite the significantly larger $\mathcal{D}$ when trained on YFCC100M, these models cannot offset the effects of the distribution shift and are outperformed by models pretrained on ImageNet in the majority of downstreams.
@@
SimCLR
$\mathbb{D}_{\text{SSL}}$ $N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
ResNet-50 ImageNet 0.128 1.31 56.9 43.0 82.3 73.9 45.4 59.9
ImageNet 0.256 2.61 61.1 45.5 84.0 76.0 47.0 64.4
ImageNet 0.640 6.54 63.7 46.8 84.2 78.8 48.3 67.2
ImageNet 1.281 13.0 64.5 46.4 84.6 79.5 48.8 68.0
[…]
ViT-B/16 ImageNet 0.128 1.31 54.2 37.2 81.8 70.3 44.6 64.3
ImageNet 0.256 2.61 61.3 39.5 83.8 77.7 47.2 69.4
ImageNet 0.640 6.54 65.5 39.9 84.6 80.4 49.1 73.1
ImageNet 1.281 13.0 66.7 39.5 83.7 81.7 50.0 73.6
@@
$\mathbb{D}_{\text{SSL}}$ $N$ $\mathcal{D}$ Accuracynormal-$\uparrow$
ResNet-50 ImageNet 0.128 1.31 58.1 40.6 81.8 76.04 45.1 63.9
ImageNet 0.256 2.61 62.9 45.3 85.0 79.8 47.6 68.7
ImageNet 0.640 6.54 65.4 48.4 86.1 81.9 49.2 71.0
ImageNet 1.281 13.0 65.9 48.8 86.6 82.6 49.5 71.9
@@
ViT-B/16 ImageNet 0.128 1.31 57.9 33.4 82.8 78.0 46.7 67.8
ImageNet 0.256 2.61 63.7 35.0 85.3 82.9 48.4 71.0
ImageNet 0.640 6.54 67.2 39.5 85.0 85.8 49.7 72.8
ImageNet 1.281 13.0 68.8 41.9 86.5 86.5 50.3 73.8

### v3

@@ Abstract
We explore the impact of training with more diverse datasets, characterized by the number of unique samples, on the performance of self-supervised learning (SSL) under a fixed computational budget. Our findings demonstrate that increasing pretraining data diversity enhances SSL performance, albeit only when the distribution distance to the downstream data is minimal. Notably, even with an exceptionally large pretraining data diversity achieved through methods like web crawling or diffusion-generated data, among other ways, the distribution shift remains a challenge. Our experiments are comprehensive with seven SSL methods using large-scale datasets such as ImageNet and YFCC100M amounting to over 200 GPU days. The code and trained models are available at https://github.com/hammoudhasan/DiversitySSL.
@@
Figure 1: Impact of Pretraining Diversity: Self-supervised learning (SSL) can be used to pretrain vision models on small datasets closely aligned to the downstream task, e.g., pets classification, hence with a small distribution shift (top, wild animals pretraining). Conversely, we could pretrain on an extensively varied dataset, with wide distribution differences (outdoor scenes, bottom). We study the role of pretraining diversity in SSL under a fixed budget, and highlight its effects in relationship to the distribution shift.
@@ 1 Introduction
Ultimately, our work provides a comprehensive analysis, from experiments worth 200 GPU days on NVIDIA A100, on the interaction between data diversity and computational resources, and their impact on the performance of SSL models.
@@ 2 Related Work
Pretraining at Scale   SSL pretraining is most effective at scale, benefiting from large datasets and great computational resources. Initial attempts at large-scale pretraining were made by combining contrastive learning and clustering [56, 7]. SEER [26] was trained on 1 billion internet images with SwAV [8] on a RegNet [49] backbone. The importance of model scaling to leverage large datasets is also shown in  [65, 28, 27], as well as the need for longer training times [65]. Additional strategies are necessary to achieve best performance [67, 17, 24]. These works focus on reaching the best representations, without considerations about training costs, thus encouraging unfair comparisons across trainings. Preliminary papers [15, 22] found that increasing data may lead to a modest boost in downstream performance when the downstream distribution changes. While they give insights on the pretraining diversity, its relation with the distribution shift under normalized computation is still not investigated.
Distribution Shift in SSL   Some studies have investigated how the pretraining data affects downstream performance on other domains. In general, the distribution shift increases when factors such as visual content and data acquisition procedure differ in two data ensembles [59]. For various pretraining datasets, preliminary works [37, 15, 28, 41] observed that the best performing representations were learned on datasets that were similar to the downstream test task. Additionally, combining datasets before pretraining or combining self-supervised features learned from various datasets did not lead to significant improvements in [15]. In [69, 52], they showcase higher generalization capabilities of SSL models compared to their supervised counterparts, for several downstream tasks in the presence of a distribution shift. In [60], they pretrained on several datasets, observing different generalization depending on the appearance of the pretraining dataset. Furthermore, initial considerations on the impact of external data on downstream tasks under distribution shift have been proposed in [61]. Some compensate the distribution shift with the scale of training datasets [30]. However, there is still a lack of a fair, computation-normalized evaluation that allows to study the effects of the distribution shift in a controlled environment.
@@ 3 Preliminaries
where $\mathcal{L}_{\text{SSL}}$ is a loss function specific to the SSL pretraining method. After pretraining, the feature extractor $f_{\theta_{f}^{*}}$ can be deployed for various downstream tasks such as image classification, object detection, or segmentation. This is typically achieved by training a task-specific head. Alternatively, the feature extractor $f_{\theta_{f}^{*}}$ can be fine-tuned or used with a $k$-nearest neighbors (kNN) classifier.
Linear Probing   There are several ways to evaluate the performance of a self-supervised learning method such as linear probing [29, 14, 10], kNN [63, 72, 9, 8], and few-shot evaluation [26, 23]. Consistently with the general protocol established in the literature [10, 32, 68, 15], we use linear probing to measure the quality of the features extracted for classification tasks. The procedure is: (1) a labeled downstream dataset, $\mathbb{D}_{\text{task}}$, consisting of image-class pairs $(\mathbf{x},\mathbf{y})\sim\mathbb{D}_{\text{task}}$, is selected for evaluation. (2) For each image $\mathbf{x}$, its representation is extracted using the pretrained feature extractor $f_{\theta_{f}^{*}}$, after which the linear classification head $t_{\theta_{t}}$, parameterized by $\theta_{t}$, is then applied to obtain $t_{\theta_{t}}(f_{\theta_{f}^{*}}(\mathbf{x}))$. (3) The linear head $t_{\theta_{t}}$ is optimized as follows:
@@
Note that only the parameters of the head $\theta_{t}$ are optimized, while the parameters $\theta^{*}_{f}$ of the feature extractor are frozen. The quality of features extracted by $f_{\theta_{f}^{*}}$ is directly inferred from the classification accuracy achieved on the test set of $\mathbb{D}_{\text{task}}$, which serves as the primary indicator of the quality of the extracted features.
@@ 4 Normalized Evaluation
[…]
×$\times 10^{3}$$\times 10^{-3}$ SimCLR B.T. BYOL SwAV VICR MoCo3 DINO
5 0.1 38.0 44.3 41.1 14.1 44.1 12.5 15.7
25 0.5 50.1 54.3 51.0 45.4 50.4 51.6 35.9
50 1.0 58.6 58.4 58.3 56.4 55.8 55.6 40.9
@@
[…]
×$\times 10^{3}$$\times 10^{-3}$ SimCLR B.T. BYOL SwAV VICR MoCo3 DINO
10 0.2 36.9 41.0 35.6 34.2 37.6 36.6 34.5
50 1.0 48.8 52.0 48.1 43.6 48.8 46.5 44.4
100 2.0 49.8 55.6 50.3 47.5 54.1 48.6 47.3
@@ 5.1 Performance on the Same Distribution
Insight (1) When the distributions of the upstream and downstream tasks are the same, i.e., $\mathbb{D}_{\text{SSL}}$ = $\mathbb{D}_{\text{task}}$, in a normalized computation setup, increasing pretraining diversity $\mathcal{D}$ proves to be an effective strategy for enhancing SSL pretraining performance.
@@ 5.2 Increasing Data Diversity
As observed in Section 5.1, if $\mathbb{D}_{\text{SSL}}=\mathbb{D}_{\text{task}}$ having a more diverse pretraining set benefits SSL methods, under a normalized computation assumption. However, to increase diversity, sampling from the same distribution $\mathbb{D}_{\text{task}}$ to extend $\mathbb{D}_{\text{SSL}}$ is not always attainable. In fact, pretraining data is often sampled from a distribution different than $\mathbb{D}_{\text{task}}$. The added samples will then introduce a distribution shift between $\mathbb{D}_{\text{SSL}}$ and $\mathbb{D}_{\text{task}}$. Then, we raise the following question: is increasing pretraining diversity still beneficial when there is a distribution shift between $\mathbb{D}_{\text{SSL}}$ and $\mathbb{D}_{\text{task}}$, i.e., $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$? We explore strategies for acquiring new data for increasing $\mathcal{D}$, namely, including existing samples, crawled internet data, and data synthetically generated. To evaluate the effects of the distribution shift in a controlled scenario, we analyze distributions closer to $\mathbb{D}_{\text{task}}$ by using a class prior (in-distribution) and without a class prior (out-of-distribution).
Figure 2: Data Collection Strategies: We analyze strategies for collecting additional data ($\mathbb{A}$), i.e., collecting more source data, crawling the web or using synthetic images. Using a class prior (top row) simulates In-distribution trainings. We also collect images without class prior (bottom row) to analyze the interactions between diversity and Out-of-distribution classes.
Setup   We use ImageNet-100 [57] as $\mathbb{D}_{\text{task}}$, and we construct multiple $\mathbb{D}_{\text{SSL}}$ to evaluate the effects of different data collection strategies and the distribution shift. We first introduce a set $\mathbb{B}$ composed of $65\times 10^{3}$ images from ImageNet-100 (50% of the dataset), which we use as a baseline for $\mathbb{D}_{\text{SSL}}$ with minimum diversity. We denote the 100 classes in ImageNet-100 as $\mathcal{T}_{\text{100}}$, and sample $\mathbb{B}$ uniformly including images from all classes. Next, we compare with pretraining on more diverse datasets as $\mathbb{D}_{\text{SSL}}$, imposing $\mathbb{D}_{\text{SSL}}=\mathbb{B}\cup\mathbb{A}$ where $\mathbb{A}$ includes $65\times 10^{3}$ images sampled with one of three strategies. To highlight the effects of the distribution shift, we include in $\mathbb{A}$ either images from In-distribution classes, i.e., selecting images from classes overlapping with $\mathcal{T}_{\text{100}}$, or images from Out-of-distribution classes, which do not overlap with $\mathcal{T}_{\text{100}}$. This is to study the effects of the distribution shift, since we do not assume access to downstream classes in real scenarios. For the Out-of-distribution samples, we define $\mathbb{A}$ as (1) random images sampled from the full ImageNet [18] dataset; (2) images crawled from Flickr, Bing, and DuckDuckGo; or (3) images generated with Stable Diffusion V2.1 [51]. We respectively call these sets $\mathbb{A}_{\text{Source}}^{\text{{{Out}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{Out}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{Out}}}}$. We also define their In-distribution counterparts as $\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{In}}}}$, respectively. Note how $\mathbb{B}\cup{\color[rgb]{0.87890625,0.4375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.87890625,0.4375,0}\mathbb{A}_{\text{Source}}^{\text{{{In}}}}}$ is the full ImageNet-100, coherently with Section 5.1. Figure 2 shows each collection strategy, for which we provide implementation details in supplementary material. Although many factors (such as the images appearance) impact the distribution shift, using a class prior with any strategy would result in closer distribution with respect to the same strategy without class priors [15]. We pretrain a ResNet-18 from scratch, with the same settings of Section 5.1, and $\mathcal{C}=50\times 10^{6}$. Note that this results in $\mathcal{D}=1.25\times 10^{-3}$ for pretraining on $\mathbb{B}$, and $\mathcal{D}=2.5\times 10^{-3}$ for pretraining on $\mathbb{B}\cup\mathbb{A}$, introducing a $\mathcal{D}$ difference of a factor of 2.
Results   We report the linear probing accuracies for pretraining on each $\mathbb{B}\cup\mathbb{A}$ as bars in Figure 3, showing the $\mathcal{C}$-normalized training on $\mathbb{B}$ as a dashed line. Surprisingly, without class priors (Out), including $\mathbb{A}_{\text{Source}}^{\text{{{Out}}}}$, $\mathbb{A}_{\text{Web}}^{\text{{{Out}}}}$, and $\mathbb{A}_{\text{Synthetic}}^{\text{{{Out}}}}$ underperforms compared to pretraining on $\mathbb{B}$ only. For instance, while SimCLR on $\mathbb{B}$ scores 72.3% accuracy, increasing diversity reduces the accuracy by 1% in all cases. This might appear to conflict with our findings in Section 5.1, however, the inclusion of Out samples leads to $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$, since we sample only classes not included in $\mathbb{D}_{\text{task}}$. We infer that, with normalized $\mathcal{C}$, increasing $\mathcal{D}$ without distribution priors may negatively impact performance. Conversely, when class priors are available (In), increasing pretraining diversity $\mathcal{D}$ improves performance compared to $\mathbb{B}$ pretraining. For instance, pretraining on $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$ performs comparably to augmenting with additional ImageNet samples ($\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$), as in the case of SimCLR where the inclusion of $\mathbb{A}_{\text{Web}}^{\text{{{In}}}}$ scores only 0.2% lower than $\mathbb{A}_{\text{Source}}^{\text{{{In}}}}$. Including $\mathbb{A}_{\text{Synthetic}}^{\text{{{In}}}}$ data also helps, although more limitedly due to the visual discrepancies between generated and real images. Ultimately, the effectiveness of diversity is linked to the distribution shift. These findings highlight the impact of the distribution shift on computationally-normalized SSL pretraining and help define evaluation practices for this task (see Section 7).
Insight (2) When the distributions of the upstream and downstream tasks differ, $\mathbb{D}_{\text{SSL}}\neq\mathbb{D}_{\text{task}}$, and in a normalized computation setup, increasing pretraining diversity $\mathcal{D}$ may harm SSL pretraining performance, reflecting the adverse effects of distribution shift.
Table 2: Performance of ImageNet and YFCC100M SSL Pretraining on Various Downstream Tasks: We train ResNet-50 and ViT-B/16 under normalized computation ($\mathcal{C}=98\times 10^{6}$) using SimCLR (left) and MoCoV3 (right) on ImageNet (IN) and YFCC100M (YFCC) with multiple $\mathcal{D}$. Despite the significantly larger $\mathcal{D}$ when trained on YFCC100M, these models cannot offset the effects of the distribution shift and are outperformed by models pretrained on ImageNet in most downstreams.
SimCLR
@@
×$\times 10^{6}$$\times 10^{-3}$ IN Cars Flow. Pets Places Food
ResNet-50 IN 0.128 1.31 56.9 43.0 82.3 73.9 45.4 59.9
IN 0.256 2.61 61.1 45.5 84.0 76.0 47.0 64.4
IN 0.640 6.54 63.7 46.8 84.2 78.8 48.3 67.2
IN 1.281 13.0 64.5 46.4 84.6 79.5 48.8 68.0
@@
ViT-B/16 IN 0.128 1.31 54.2 37.2 81.8 70.3 44.6 64.3
IN 0.256 2.61 61.3 39.5 83.8 77.7 47.2 69.4
IN 0.640 6.54 65.5 39.9 84.6 80.4 49.1 73.1
IN 1.281 13.0 66.7 39.5 83.7 81.7 50.0 73.6
@@
×$\times 10^{6}$$\times 10^{-3}$ IN Cars Flow. Pets Places Food
ResNet-50 IN 0.128 1.31 58.1 40.6 81.8 76.04 45.1 63.9
IN 0.256 2.61 62.9 45.3 85.0 79.8 47.6 68.7
IN 0.640 6.54 65.4 48.4 86.1 81.9 49.2 71.0
IN 1.281 13.0 65.9 48.8 86.6 82.6 49.5 71.9
[…]
ViT-B/16 IN 0.128 1.31 57.9 33.4 82.8 78.0 46.7 67.8
IN 0.256 2.61 63.7 35.0 85.3 82.9 48.4 71.0
IN 0.640 6.54 67.2 39.5 85.0 85.8 49.7 72.8
IN 1.281 13.0 68.8 41.9 86.5 86.5 50.3 73.8